# Identification of three new ULXs in NGC 4631 and NGC 1097: evidence for stellar-mass black holes

### v1

@@ 4.0.3 Optical counterparts of ULX X-6 and X-7
The optical counterparts of ULXs X-6 and X-7 appear faint in both filters, consistent with the magnitudes of typical ULX optical counterparts. Since multiple possible counterparts are detected for both sources, it is not possible to place reliable constraints on their optical emission. However, the CMDs allow us to derive some constraints on their ages and masses. However, all three candidate counterparts of X-6 are located at positions on the CMD that are clearly distinct from the field stars in the region (see Figure 9), suggesting that they might belong to a different stellar population or might have followed a different evolutionary path. On the other hand, the source X6a, which is relatively brighter than the other counterparts, is distinct from the isochrones on the CMD. For the candidate counterpart X6a, the exceptionally blue color may be affected by observational limitations. Although identical photometric procedures were applied to all sources, the HST data are limited and non-simultaneous, leaving the optical variability of X6a unknown. If X6a is the true counterpart, its optical emission may be dominated by the accretion disk rather than the donor star, which would produce a displaced position in the CMD. Deeper and simultaneous optical and X-ray observations are required to establish the true properties of X6a.
@@
Based on the CMD, X-6b and X-6c, as well as X-7a and X-7b, have estimated ages of 30–40 Myr and masses in the range 7–9 M⊙, consistent with an HMXB classification. In addition, due to the lack of continuous optical monitoring and simultaneous optical and X-ray observations, the variability of the candidate counterparts remains uncertain. Consequently, we could not derive the FX/Foptic ratio for X-6 and X-7, which serves to constrain the nature of sources through their X-ray-to-optical flux comparison (Maccacaro et al., 1988; Stocke et al., 1991).
@@ 4.1 NGC 1097 ULX-3
NGC 1097 ULX-3 is analyzed in detail for the first time in this study and identified as an ULX based on its X-ray spectral and temporal properties, with additional observations in the optical and infrared bands. The ULX-3 was re-detected in the most recent five Chandra observations between 21 May and 31 May 2025 consistently with a stable LX (6–7) 1039 erg s-1, showing no significant variability on short timescales. The XMM-Newton observation performed in 2023 showed a comparable luminosity level of erg s-1. The long-term monitoring with Swift/XRT between 2 August 2007 and 6 June 2025 reveals significant variability. Over this 18-year period, the LX of the ULX-3 varied between and erg s-1. Such long-term flux variation suggests that ULX-3 is a highly variable source, likely driven by significant changes in the accretion rate, accretion state transitions, or geometric effects such as variable absorption or precession of the accretion disk. Overall, these results suggest that ULX-3 remains relatively stable on short timescales, as indicated by the recent Chandra and XMM-Newton observations, but exhibits obvious variability over long timescales spanning years to decades, as revealed by the Swift/XRT monitoring.
@@
The X-ray spectrum of the ULX-3 is best described by a two-component model consisting of a soft thermal blackbody and a hard power-law (power-law + bbody). The soft thermal component, with a temperature of keV, is consistent with the soft excess observed in many ULXs (Gladstone et al., 2009; Sutton et al., 2013), and is typically attributed to emission from an optically thick wind or outflow launched from a supercritical accretion disk (Poutanen et al., 2007; Middleton et al., 2015). This component may correspond to the photosphere of the wind or to the outer, geometrically thick regions of the accretion flow. Such hard spectral slopes are commonly associated with the so-called HUL in ULXs (Gladstone et al., 2009; Sutton et al., 2013; Heida et al., 2014), where the inner regions of the accretion flow may be partially visible through the wind funnel, allowing Comptonized emission from a hot corona or the innermost disk to dominate at higher energies. The combination of a low-temperature thermal component and a hard power-law tail supports a super-Eddington accretion scenario for the ULX-3, likely involving a stellar-mass compact object surrounded by a geometrically and optically thick accretion structure.(Pinto et al., 2021).
@@
To investigate spectral state transitions, we constructed the hardness evolution and hardness intensity diagram shown in the b and c panels of Figure 6. In high count rates, the source spectrum becomes softer, consistent with a thermal-disk dominated (high/soft) state, whereas at low count rates the spectrum hardens, indicative of a corona or Comptonization-dominated (low/hard) state. Such a correlation is a signature of accretion-powered variability, linking luminosity changes to transitions between spectral states. In Galactic XRBs, such transitions are well-characterized and often correspond to changes between low/hard and high/soft states driven by variations in mass accretion rate (Remillard and McClintock, 2006; Done et al., 2007). In ULXs, however, the spectral state transitions are more complex due to their potential super-Eddington luminosities and geometric beaming effects (King, 2009). The observed transitions in this source may reflect different accretion modes onto a compact object, which could be either a stellar-mass BH or an NS (Poutanen et al., 2007; Bachetti et al., 2014; Middleton et al., 2015; Walton et al., 2018). Optically thick radiatively driven outflows could obscure the inner hot regions of the disk, leading to softer spectra when observed at moderate inclination angles. In case of harder spectra at even higher flux levels, the changes in the wind geometry, collimation, or the re-emergence of the innermost accretion flow could indicate indicate a transition in the accretion regime (Kawashima et al., 2012). Overall, the spectral state transitions observed in NGC 1097 ULX-3 reinforce the emerging picture that ULXs form a heterogeneous population, displaying diverse accretion behaviors that span the regime between Galactic XRBs and super-Eddington accretion, implying that ULX-3 could be powered either by a stellar-mass black hole or an NS.
4.1.1 Optical and near-infrared
@@
We identified a unique optical and NIR counterpart within the astrometric error radius of the ULX-3. The optical counterpart is fainter than typical ULX counterparts, with F555W mag ( mag). In the F814W filter, the variability of the source is quantified as the difference between the faintest and brightest magnitudes measured across three epochs (2004–2019), yielding mag, suggesting intrinsic variability in the optical emission of ULX-3. The variable X-ray emission is likely linked to the observed optical variability, suggesting that accretion processes dominate the optical emission.
@@
In contrast, the NIR observations (2023-12-27-2024-09-29) reveal that the source remains almost constant in magnitudes. Furthermore, we can exclude the scenario of a variable jet contribution to the NIR emission, since the infrared emission remains constant. Using simultaneous JWST/NIRCam filters, we constructed the spectral energy distribution (SED) of the ULX counterpart and fitted it with a blackbody model with a temperature of 3300 K (see Figure 11. The derived effective temperature is too high to be attributed to circumbinary material/disk or dust (Dudik et al., 2016; Lau et al., 2017, 2019; Allak, 2023), and instead falls within the expected range for RSGs. Adopting a blackbody temperature of 3300 K, we used the relation with the NIRCam photometry at a distance of 16.8 Mpc to derive an RSG radius of . The absolute magnitude of the counterpart is mag. In the NIR CMD, this value lies on the evolutionary tracks of stars with ages of 7–12 Myr, indicating that its CMD position is consistent with the RSG evolutionary branch.
@@ 5 Conclusions
In this work, we reported the identification and characterization of three new ULXs in NGC 4631 and NGC 1097. Our main findings can be summarized as follows:
NGC 4631 X-6: The source shows LX variability of about an order of magnitude and recurrent spectral variability. Spectral analyses support its classification as a stellar-mass BH candidate accreting in a super-Eddington regime. The – relation follows the trend expected for a standard thin disk, supporting the interpretation of a stellar-mass BH accretor.
NGC 4631 X-7: This transient ULX displays long-term LX variability and disk-dominated spectra with inner disk temperatures of 0.5–1 keV. Its photometric properties are consistent with a donor star of 7–9 M and an age of 30–40 Myr, supporting an HMXB scenario. Moreover, trend expected for standard thin disks, providing evidence for a slim disk accretion state. Finally, both ULXs X-6 and X-7, our analysis indicates that their properties are consistent with normal HMXBs hosting BHs with masses of 20 M.
NGC 1097 ULX-3: The source exhibits strong long-term variability, with LX variability up to a factor of 30. Spectral state transitions indicate a stellar-mass BH or NS accretor in a super-Eddington state. We identified a unique optical and NIR counterpart. While the optical emission is highly variable, the NIR counterpart remains nearly constant. For the optical counterpart of ULX-3, the NIR SED is consistent with a blackbody temperature of K and a stellar radius of 200 R⊙, providing strong evidence that the source is a RSG donor star.
For the three ULXs in the two host galaxies, additional multi-wavelength observations will be crucial for constraining their environments, donor stars, and accretion processes, ultimately leading to a clearer understanding of their nature.
@@ Acknowledgements.
- New transient ULX candidate in NGC 4254: evidence of circumbinary disc?. MNRAS 526 (4), pp. 5765–5776. External Links: Document, 2310.09421, ADS entry Cited by: §1, §1, §3.4, §4.1.1.
@@
- Spitzer IRAC Observations of IR Excess in Holmberg IX X-1: A Circumbinary Disk or a Variable Jet?. ApJ 831 (1), pp. 88. External Links: Document, 1608.05348, ADS entry Cited by: §1, §4.1.1.
@@
- First Detection of Mid-infrared Variability from an Ultraluminous X-Ray Source Holmberg II X-1. ApJ 838 (2), pp. L17. External Links: Document, 1703.03802, ADS entry Cited by: §4.1.1.
- Uncovering Red and Dusty Ultraluminous X-Ray Sources with Spitzer. ApJ 878 (1), pp. 71. External Links: Document, 1904.09852, ADS entry Cited by: §1, §4.1.1.
@@
Appendix A X-ray, optical and NIR observations of the ULXs
@@
| CH3 | 30937 | ACIS-S | 9.94 | 2025-05-21 |
| CH4 | 26485 | ACIS-S | 9.94 | 2025-05-21 |
| CH5 | 26486 | ACIS-S | 9.94 | 2025-05-21 |
| CH6 | 30948 | ACIS-S | 14.71 | 2025-05-31 |
| CH7 | 25782 | ACIS-S | 14.71 | 2025-05-31 |
@@
| F150W | jw03707 | NIRCam/IMAGE | 0.22 | 2023-12-27 |
| F150W | jw04793 | NIRCam/IMAGE | 0.86 | 2024-09-29 |

### v2

@@ 4.1.3 Optical counterparts of ULXs X-6 and X-7
The optical counterparts of ULXs X-6 and X-7 appear faint in both filters, consistent with the typical magnitudes observed for ULX optical counterparts. Because multiple possible counterparts are detected for both sources, we cannot place reliable constraints on their optical emission. Nevertheless, the CMDs allow us to derive some constraints on their ages and masses. However, all three candidate counterparts of X-6 occupy positions on the CMD that are clearly distinct from the field stars in the surrounding region (Figure 9), suggesting that they may belong to a different stellar population or have followed a different evolutionary path. The source X6a, which is relatively brighter than the other counterparts, is distinct from the isochrones on the CMD. The exceptionally blue color of X6a may be affected by observational limitations. Although we applied identical photometric procedures to all sources, the HST data are limited and nonsimultaneous, leaving the optical variability of X6a unknown. If X6a is the true counterpart, its optical emission may be dominated by the accretion disk rather than by the donor star, which would produce a displaced position in the CMD. Deeper and simultaneous optical and X-ray observations are required to establish the true properties of X6a.
@@
Based on the CMDs, X-6b, X-6c, X-7a, and X-7b have estimated ages of 30–40 Myr and masses in the range 7–9 M⊙, consistent with an HMXB classification. Because continuous optical monitoring and simultaneous optical and X-ray observations are unavailable, the variability of these candidate counterparts remains uncertain. Consequently, we could not derive the FX/Foptic ratio for X-6 and X-7, which serves to constrain the nature of the sources through their X-ray-to-optical flux comparison (Maccacaro et al., 1988; Stocke et al., 1991).
@@ 4.2 NGC 1097 ULX-3
In this study, we analyzed NGC 1097 ULX-3 in detail for the first time and identify it as an ULX based on its X-ray spectral and temporal properties, with additional observations in the optical and infrared bands. We detected ULX-3 in the five most recent Chandra observations between 21 May and 31 May 2025, with a stable LX (6–7) 1039 erg s-1 and no significant variability on short timescales. The XMM-Newton observation performed in 2023 shows a comparable luminosity level of erg s-1. Long-term monitoring with Swift/XRT between 2 August 2007 and 6 June 2025 reveals significant variability. During this 18-year period, LX of ULX-3 varied between and erg s-1. Such long-term flux variation suggests that ULX-3 is a highly variable source, likely driven by significant changes in the accretion rate, accretion state transitions, or geometric effects such as variable absorption or accretion-disk precession. Overall, these results suggest that ULX-3 remains relatively stable on short timescales, as indicated by the recent Chandra and XMM-Newton observations, but exhibits obvious variability over long timescales spanning years to decades, as revealed by the Swift/XRT monitoring.
@@
The X-ray spectrum of ULX-3 is best described by a two-component model consisting of a soft thermal blackbody and a hard power-law component (power-law plus bbody). The soft thermal component, with a temperature of keV, is consistent with the soft excess observed in many ULXs (Gladstone et al., 2009; Sutton et al., 2013) and is typically attributed to emission from an optically thick wind or outflow launched from a supercritical accretion disk (Poutanen et al., 2007; Middleton et al., 2015). This component may correspond to the wind photosphere or to the outer, geometrically thick regions of the accretion flow. Such hard spectral slopes are commonly associated with the so-called HUL state in ULXs (Gladstone et al., 2009; Sutton et al., 2013; Heida et al., 2014), in which the inner regions of the accretion flow may be partially visible through the wind funnel, allowing Comptonized emission from a hot corona or the innermost disk to dominate at higher energies. The combination of a low-temperature thermal component and a hard power-law tail supports a super-Eddington accretion scenario for ULX-3, likely involving a stellar-mass compact object surrounded by a geometrically and optically thick accretion structure (Pinto et al., 2021).
@@
To investigate spectral state transitions, we constructed hardness evolution and hardness intensity diagrams (Figure 6, panels b and c). At high count rates, the source spectrum becomes softer, consistent with a thermal disk-dominated (high-soft) state, whereas at low count rates, the spectrum hardens, indicative of a corona- or Comptonization-dominated (low-hard) state. Such a correlation is a signature of accretion-powered variability, linking luminosity changes to transitions between spectral states. In Galactic XRBs, such transitions are well characterized and often correspond to changes between low-hard and high-soft states driven by variations in the mass accretion rate (Remillard and McClintock, 2006; Done et al., 2007). In ULXs, however, spectral state transitions are more complex due to their potential super-Eddington luminosities and geometric beaming effects (King, 2009). The observed transitions in this source may reflect different accretion modes onto a compact object, which could be either a stellar-mass BH or an NS (Poutanen et al., 2007; Bachetti et al., 2014; Middleton et al., 2015; Walton et al., 2018). Optically thick, radiatively driven outflows could obscure the inner hot regions of the disk, leading to softer spectra when observed at moderate inclination angles. In the case of harder spectra at even higher flux levels, changes in wind geometry, collimation, or the reemergence of the innermost accretion flow could indicate a transition in the accretion regime (Kawashima et al., 2012). Overall, the spectral state transitions observed in NGC 1097 ULX-3 reinforce the emerging picture that ULXs form a heterogeneous population, displaying diverse accretion behaviors that span the regime between Galactic XRBs and super-Eddington accretion, implying that ULX-3 could be powered either by a stellar-mass BH or an NS.
@@
We identified a unique optical and NIR counterpart within the astrometric error radius of ULX-3. The optical counterpart is fainter than typical ULX counterparts, with F555W mag ( mag). In the F814W filter, the variability of the source is quantified as the difference between the faintest and brightest magnitudes measured over three epochs (2004–2019), yielding mag and suggesting intrinsic variability in the optical emission of ULX-3. The variable X-ray emission is likely linked to the observed optical variability, suggesting that accretion processes dominate the optical emission.
@@
In contrast, the NIR observations obtained between 27 December 2023 and 29 September 2024 reveal that the source remains nearly constant in magnitude. Furthermore, we can exclude a variable jet contribution to the NIR emission, since the infrared emission remains constant. Using simultaneous JWST/NIRCam filters, we constructed the spectral energy distribution (SED) of the ULX counterpart and fit it with a blackbody model with a temperature of 3300 K (see Figure 11). The effective temperature derived is too high to be attributed to circumbinary material or dust (Dudik et al., 2016; Lau et al., 2017, 2019; Allak, 2023) and instead falls within the expected range for RSGs. Adopting a blackbody temperature of 3300 K, we used the relation , together with NIRCam photometry at a distance of 16.8 Mpc, to derive an RSG radius of . The absolute magnitude of the counterpart is mag. In the NIR CMD, this value lies on the evolutionary tracks of stars with ages of 7–12 Myr, indicating that its CMD position is consistent with the RSG evolutionary branch.
@@ 5 Conclusions
In this work, we report the identification and characterization of three new ULXs in NGC 4631 and NGC 1097. Our main findings can be summarized as follows.
NGC 4631 X-6 exhibits LX variability of about an order of magnitude and recurrent spectral variability. Spectral analyses support its classification as a stellar-mass BH candidate accreting in a super-Eddington regime. The – relation follows the trend expected for a standard thin disk, supporting the interpretation of X-6 as a stellar-mass BH accretor.
NGC 4631 X-7, a transient ULX, displays long-term LX variability and disk-dominated spectra with inner disk temperatures of 0.5–1 keV. Its photometric properties are consistent with a donor star of 7–9 M and an age of 30–40 Myr, supporting an HMXB scenario. Moreover, the trend expected for standard thin disks provides evidence for a slim disk accretion state. Finally, our analysis indicates that both X-6 and X-7 have properties consistent with normal HMXBs hosting BHs with masses of 20 M.
NGC 1097 ULX-3 exhibits strong long-term variability, with LX changes up to a factor of 30. Spectral state transitions indicate that ULX-3 is a stellar-mass BH or NS accretor in a super-Eddington state. We identify a unique optical and NIR counterpart to ULX-3. Although the optical emission is highly variable, the NIR counterpart remains nearly constant. For the optical counterpart of ULX-3, the NIR SED is consistent with a blackbody temperature of K and a stellar radius of 200 R⊙, providing strong evidence that the source is an RSG donor star.
For the three ULXs in the two host galaxies, additional multiwavelength observations will be crucial to constrain their environments, donor stars, and accretion processes, ultimately leading to a clearer understanding of their nature.
@@ Acknowledgements.
- New transient ULX candidate in NGC 4254: evidence of circumbinary disc?. MNRAS 526 (4), pp. 5765–5776. External Links: Document, 2310.09421, ADS entry Cited by: §1, §1, §3.4, §4.2.
@@
- Spitzer IRAC Observations of IR Excess in Holmberg IX X-1: A Circumbinary Disk or a Variable Jet?. ApJ 831 (1), pp. 88. External Links: Document, 1608.05348, ADS entry Cited by: §1, §4.2.
@@
- First Detection of Mid-infrared Variability from an Ultraluminous X-Ray Source Holmberg II X-1. ApJ 838 (2), pp. L17. External Links: Document, 1703.03802, ADS entry Cited by: §4.2.
- Uncovering Red and Dusty Ultraluminous X-Ray Sources with Spitzer. ApJ 878 (1), pp. 71. External Links: Document, 1904.09852, ADS entry Cited by: §1, §4.2.
@@
Appendix A X-ray, optical, and NIR observations of the ULXs
@@
| CH3 | 30142 | ACIS-S | 9.94 | 2025-05-21 |
| CH4 | 30937 | ACIS-S | 9.94 | 2025-05-21 |
| CH5 | 30938 | ACIS-S | 9.94 | 2025-05-21 |
| CH6 | 29637 | ACIS-S | 14.71 | 2025-05-31 |
| CH7 | 30948 | ACIS-S | 14.71 | 2025-05-31 |
@@
| F115W | jw03707 | NIRCam/IMAGE | 0.22 | 2023-12-27 |
| F115W | jw04793 | NIRCam/IMAGE | 0.86 | 2024-09-29 |